OpenStudy (anonymous):

Meg opened a savings account with $44.50. She puts in $22.25 in the account each week when she gets paid. She is saving to purchase a blu-ray player that costs $178.00. How many weeks will it take her to save enough money? A. 6 weeks B. 7 weeks C. 8 weeks D. 9 weeks

OpenStudy (anonymous):

A. if you take 178.00- 44.50 you that the answer from that and divide it by 22.25 and get your answer.

OpenStudy (anonymous):

answer = A.

OpenStudy (anonymous):

one week \(44.50+22.25\) two weeks \(44.50+22.25\times 2\) three weeks \(44.50+22.25\times 3\) n weeks \[44.50+22.25\times n\]

OpenStudy (anonymous):

^ is a more simple way to put it.